OpenStudy (anonymous):

Need some help with this please... f(x)=(squareroot)2-2x x=-4,4 Need to figure out what f(-4)=

OpenStudy (johnweldon1993):

I'm confused as to what the x = -4,4 is? but for f(-4) all you should have to do...is plug in (-4) for every x you see f(x) = √2-2(-4) = √2-(-8) = √10 definietly not 100% on this because like I said idk what the x = -4,4 means here

OpenStudy (anonymous):

yeah the -4, 4 is the #'s that we are to substitute

OpenStudy (anonymous):

I see about plugging in the #'s for x. Thank you John :)

OpenStudy (johnweldon1993):

lol *even though I didn't help much lol* you are very welcome :)